OpenStudy (anonymous):

Let f=R->R be such that f(1)=3 and f '(1)=6 , then lim(x tends to zero) [(f(1+x)/f(1)]^(1/x) equals to ----------

OpenStudy (anonymous):

Let f=R->R be such that f(1)=3 and f '(1)=6 , then \[\lim_{x \rightarrow 0}[(f(1+x)/f(1)]^{1/x}\] equals to ----------

OpenStudy (anonymous):

infinity, I guess :) since the power will be infinity, anything to the power of infinity is infinity :)

OpenStudy (anonymous):

or you can solve it using e and ln too

OpenStudy (anonymous):

wait....that's not the case here, what's the function?

OpenStudy (nowhereman):

you don't need the function Just write it as \[\exp\left(\frac {\ln \frac{ f(1+x)}{f(1)}}{x}\right)\] and you can solve it with l'hopital.

OpenStudy (anonymous):

Could you please explain how to come to that step?

OpenStudy (nikvist):

limes doesn't exist, because \[\lim_{x\rightarrow 0-}[f(1+x)/f(1)]^{1/x}\neq\lim_{x\rightarrow 0+}[f(1+x)/f(1)]^{1/x}\]

OpenStudy (nikvist):

wrong, my last post is not true

OpenStudy (nowhereman):

I just put exp(ln(.)) around the function and used logarithm rules.

OpenStudy (anonymous):

Okay, I will certainly look at it, though I have to leave again. If I can get a chance to post tonight, I will.

OpenStudy (anonymous):

That alright?

OpenStudy (anonymous):

Ok, al right

OpenStudy (nowhereman):

What is it you don't understand?

OpenStudy (anonymous):

Its ok, I understand what you said

OpenStudy (anonymous):

I have something else to ask

OpenStudy (anonymous):

I just wish to ask, what are the things to consider while solving problems on limit and continuity

OpenStudy (nowhereman):

I'm sorry, but I don't know a general answer to that question. Analysis is a huge field.

OpenStudy (anonymous):

No no, you don't have to go to that. I am in high school

OpenStudy (anonymous):

AYCC is a homework tutoring marketplace. If you're a student seeking smart solutions to your difficult homework assignments or someone who’d love to pass on your pearls of wisdom for a specific topic or a question and earn additional income, join www.aceyourcollegeclasses.com

OpenStudy (anonymous):

iam, I can respond now if you're around.

OpenStudy (anonymous):

Yes I am online

OpenStudy (anonymous):

\[y=\left( \frac{f(x+1)}{f(1)} \right)^{1/x} \rightarrow \log y = \log \left( \frac{f(x+1)}{f(1)} \right)^{1/x}\]That is\[\log y =\frac{1}{x}\log \left( \frac{f(x+1)}{f(1)} \right)=\frac{\log \left( \frac{f(x+1)}{f(1)} \right)}{x}\]

OpenStudy (anonymous):

Then\[\lim_{x->\infty}\log y=\lim_{x->\infty}\frac{\log \left( \frac{f(x+1)}{f(1)} \right)}{x}\]

OpenStudy (anonymous):

Indeterminate, 0/0 (change that limit to 0, not infinity).

OpenStudy (anonymous):

Use L'Hopital's rule. Then\[\lim_{x \rightarrow 0}\frac{\log \frac{f(x+1)}{f(1)}}{x}=\frac{\lim_{x \rightarrow 0}\log \frac{f(x+1)}{f(1)}}{\lim_{x \rightarrow 0}x}\]\[=\frac{\lim_{x \rightarrow 0}\frac{\left( \frac{f'(x+1)}{f(1)} \right)}{\left( \frac{f(x+1)}{f(1)} \right)}}{\lim_{x \rightarrow 0}1} =\frac{\lim_{x \rightarrow 0}f'(x+1)}{\lim_{x \rightarrow 0}f(x+1)}\]

OpenStudy (anonymous):

No need to reply to this message. Just informing that I have received the notes

OpenStudy (anonymous):

\[=\frac{f'(1)}{f(1)}=\frac{6}{3}=2\]

OpenStudy (anonymous):

I have got that. I can handle the rest.

OpenStudy (anonymous):

That is\[\lim_{x \rightarrow 0}\log y = 2 \rightarrow \lim_{x \rightarrow 0}y=e^2\]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

You are really a good person. In this world no one does so much for any other without any return

OpenStudy (anonymous):

Sometimes they do.

OpenStudy (anonymous):

This site has a good summary of limit laws: http://www.math.oregonstate.edu/home/programs/undergrad/CalculusQuestStudyGuides/SandS/lHopital/limit_laws.html

OpenStudy (anonymous):

Ok let me study the whole thing, I will let you know when I need any help. Thank you sooooooooo much

OpenStudy (anonymous):

Is that your university?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Anyhow thanks

OpenStudy (anonymous):

np. good luck

OpenStudy (anonymous):

I don't want to be nosy about you, as it may offend you

OpenStudy (anonymous):

After all I am receiving something GREAT from you

OpenStudy (anonymous):

It's okay. Just do well in STEP.

OpenStudy (anonymous):

It will be a gift from me to you

OpenStudy (anonymous):

I will do good, and thats a promise

OpenStudy (anonymous):

Well, it's worth it then :)

OpenStudy (anonymous):

Ok bye now

OpenStudy (anonymous):

See you.